OpenStudy (anonymous):

Can anyone tell me the answer to this. Predict a particular compound has the empirical formula CH2O. Its molar mass is 240.0 g/mol. Predict the molecular formula for this compound.

OpenStudy (anonymous):

First find the molecular mass of CH2O= 12+2(1)+16=30g/mol Secondly, find the dividend of the total molecular mass= 240/30 = 8 Therefore the molecular formula is 8(CH2O)= C8H16O8.